# A Framework for Quantum Simulations of Energy-Loss and Hadronization in Non-Abelian Gauge Theories: SU(2) Lattice Gauge Theory in 1+1D

### v2

@@ I Introduction
Extensive experimental campaigns in high-energy heavy-ion collisions are designed to reveal the structure and dynamics of matter under extreme conditions. The remarkable discovery of a near-ideal, low-viscosity quantum liquid of quarks and gluons produced in such collisions [PHENIX:2004vcz, BRAHMS:2004adc, PHOBOS:2004zne, STAR:2005gfr] has driven a program to measure and interpret its properties. An important line of investigation is to quantify properties of transport in the liquid phase, where the identification of signatures and probes present a challenge. One class of probes is the production, motion and measurement of heavy quarks, including the production rates of quarkonia, and processes such as dissociation and recombination. Produced in hard collisions at early times, heavy quarks witness the entire evolution of the exotic matter and its transition to hadronic final states. Therefore, the measured relative yields of heavy-quark systems reveal key attributes of the medium. An excellent review of the state-of-the-art in this area can be found in Ref. [He_2023]. An important element in extracting properties of this exotic matter is understanding how heavy quarks move and interact within the evolving system, including how they lose and gain energy as they move. From a theoretical standpoint, first-principles simulations of these collisions present a monumental challenge. This challenge is not limited to evolution through exotic matter. Even simulations of high-energy probes impacting cold dense matter, such as a nucleus, or nuclear matter, suffer the same difficulties. While processes occurring over short-distances can be reliably determined in perturbation theory, the subsequent evolution (cascade) to longer length scales involves fragmentation and hadronization, both of which are intrinsically non-perturbative. Tuned to experimental data, modeling such processes is impressively successful, e.g., Geant4 [GEANT4:2002zbu, Muskinja:2020ots, basaglia2024geant4gamechangerhigh, Pythia8], but remains incomplete in describing fundamentally quantum observables, e.g., Refs. [Bauer:2019qxa, Klco:2021lap, Bauer:2022hpo, Gray:2022fou, Bauer:2023qgm, DiMeglio:2023nsa]. While classical Euclidean-space simulations of quantum chromodynamics (QCD) encounter a sign problem in summing complex phases in the time-evolution, real-time evolution, starting from a specified initial state, is known to be efficient for quantum computers.
@@
Constrained by available noisy intermediate-scale quantum (NISQ) [Preskill:2018jim] era quantum hardware, present-day quantum simulations of quantum field theories are focused on modest-sized, low-dimensional systems [Martinez:2016yna, Klco:2018kyo, Kokail:2018eiw, Lu:2018pjk, Klco:2019evd, Mil:2019pbt, Yang:2020yer, Ciavarella:2021nmj, Bauer:2021gup, Atas:2021ext, ARahman:2021ktn, Mazzola:2021hma, deJong:2021wsd, Gong:2021bcp, Zhou:2021kdl, Riechert:2021ink, Ciavarella:2021lel, Nguyen:2021hyk, Su:2022glk, Illa:2022jqb, Mildenberger:2022jqr, ARahman:2022tkr, Asaduzzaman:2022bpi, Farrell:2022wyt, Atas:2022dqm, Farrell:2022vyh, Mueller:2022xbg, Pomarico:2023png, Charles:2023zbl, Zhang:2023hzr, Ciavarella:2023mfc, Farrell:2023fgd, Meth:2023wzd, Borzenkova:2023xaf, Schuster:2023klj, Angelides:2023noe, Farrell:2024fit, Kavaki:2024ijd, Davoudi:2024wyv, Turro:2024pxu, Ciavarella:2024fzw, Gyawali:2024hrz, Gonzalez-Cuadra:2024xul, Zemlevskiy:2024vxt, Crippa:2024hso, Ciavarella:2024lsp, Than:2024zaj, Turro:2025sec, Alexandrou:2025vaj, Farrell:2025nkx, Davoudi:2025rdv, Schuhmacher:2025ehh, Chernyshev:2025lil, Chen:2025zeh, Cobos:2025krn, Saner:2025nrq, Than:2025gso], with the Schwinger model [Schwinger:1962tp] (quantum electrodynamics (QED) in 1+1D) receiving significant attention. Like QCD, the Schwinger model is a confining gauge theory, and remains a useful test-bed for exploring new algorithms and workflows and gaining improved understanding of infinite-volume and continuum limits. Our present work is related to simulating transport, in a general sense, in non-equilibrium non-Abelian systems, and is on the path for robust simulations of jet physics and hadronization in QCD. There is a growing foundation for such simulations, e.g., Refs. [Kuhn:2015zqa, Kasper:2015cca, Sala:2018dui, Surace:2019dtp, Magnifico:2019kyj, Alexandru:2019nsa, Li:2020uhl, Bauer:2021gup, Ji:2020kjk, Barata:2021yri, Li:2021zaw, Yao:2022eqm, Barata:2022wim, Carena:2022kpg, Florio:2023dke, Barata:2023clv, Florio:2024aix, Alexandrou:2025vaj, Barata:2025hgx], along with progress in string breaking [Hebenstreit:2013baa, Lee:2023urk, Belyansky:2023rgh, Barata:2023jgd, De:2024smi, Surace:2024bht, Ciavarella:2024lsp, Gonzalez-Cuadra:2024xul, grieninger2025st, Borla:2025gfs, Xu:2025abo, Tian:2025mbv, Cataldi:2025cyo], and of the associated entanglement evolution [Chai:2023qpq, Papaefstathiou:2024zsu]. Simulations of SU(2) and SU(3) theories including quarks and anti-quarks in 1+1D, with the gluon fields constrained by Gauss’s law (either as dynamical fields or through long-range interactions), are now being performed using quantum computers [Klco:2019evd, Ciavarella:2021nmj, Atas:2021ext, Farrell:2022wyt, Farrell:2022vyh, Atas:2022dqm, Than:2024zaj, Chernyshev:2025lil]. With the Jordan-Wigner mapping of quarks, each spatial lattice site requires $2n_{f}N_{c}$ qubits for $n_{f}$ flavors of light quarks with $N_{c}$ colors. The inclusion of electroweak processes [Farrell:2022vyh, Chernyshev:2025lil] (using the low-energy effective field theory of the Standard Model) requires additional $4n_{g}$ qubits encoding the charged-leptons and neutrinos, where $n_{g}$ is the number of generations included in the simulation. Progress has also been made toward simulations using qudit systems, e.g., Refs. [Gustafson:2021qbt, Meth:2023wzd, Zache:2023cfj, Gustafson:2023kvd, Calajo:2024qrc, Illa:2024kmf, Gustafson:2024kym, Crane:2024tlj, Kurkcuoglu:2024cfv, Gaz:2025ptu, Jiang:2025ufg, Huie:2025yzn, Kurkcuoglu:2025gik, Perez:2025cxl], which have significant advantages in circuit depth and execution times. Additionally, there have been a number of important quantum and classical algorithm developments for wavepacket preparation for scattering [Chai:2023qpq, Farrell:2024fit, Farrell:2024mgu, Davoudi:2024wyv, Zemlevskiy:2024vxt, Turco:2025jot, Farrell:2025nkx, Chai:2025qhf, Davoudi:2025rdv], and for systematically-improvable truncation schemes [Li:2024lrl].
@@
Building on our previous work that established a framework for quantum simulations of energy-loss and fragmentation in systems of electrically-charged particles in the context of Abelian lattice gauge theory (LGT) [Farrell:2024mgu], we now address energy loss in non-Abelian lattice gauge theories. [^1] This should be considered to be a specific case of quantum thermodynamics of a non-equilibrium process [YungerHalpern:2019mvj, Davoudi2024a], in the context of a non-Abelian lattice gauge theory. In particular, we establish a framework to simulate the energy loss and collisions of particles in an SU(2) non-Abelian lattice gauge theory in 1+1D using quantum computers. Static heavy quark fields are mapped to qubits in order to simulate hadron motion and collisions on the lattice (as well as to establish regions of dense matter). To prepare ground-state wavefunctions in the presence of a heavy quark, SC-ADAPT-VQE [Grimsley_2019, Farrell:2023fgd] is combined with Domain Decomposition (DDec), an algorithm used extensively in traditional lattice QCD calculations [10.1093/oso/9780198501787.001.0001, saad2003iterative, Luscher:2003qa, Heybrock2014LatticeQW, frommer2014DD, Ce:2016idq], leveraging the heterogeneity of charge sectors across different regions on the lattice. The heavy quarks are translated across the lattice via discrete fermionic-SWAP operators, introducing additional finite lattice-spacing artifacts that vanish in the continuum limit. The states are then time-evolved to allow the light quarks to respond dynamically to the heavy quark motion. Scalable quantum circuits [Farrell:2023fgd] are developed for state preparation, discrete translation, and time evolution, enabling measurements of the time dependence of local non-Abelian charge fluctuations, entanglement, and the energy in the light quarks and gluons. Complete simulations (state preparation, translation, time evolution and measurement, with associated mitigation circuits) of a lattice with $L=3$ spatial sites are performed, requiring a total of 18 qubits, using IBM’s heron-processor quantum computers, with a transpiled CNOT depth of $\sim 400$. With multiple error-mitigation strategies applied, the results we obtain from the quantum simulations are in good agreement with corresponding classical simulations.
@@ II The SU(2) Lattice with Heavy Quarks
Contributions to the Hamiltonian from the chromo-electric fields are determined by combinations of the non-Abelian SU(2) charge operators [Atas:2021ext],
@@
where $\lfloor y\rfloor$ denotes the floor of $y$. With OBCs, and the requirement that the chromo-electric field outside of the lattice ($x<0$ and x>L1x>L-1) has to vanish, the summation is restricted to the electric-field boundaries in the interior of the lattice, $b\in\{0,1,\ldots,2L-2\}$. Color neutrality of the lattice is enforced by including a large energy penalty term in the Hamiltonian for non-zero $E_{b}^{(a)}$ at b=2L1b=2L-1 [Farrell:2022wyt]. The chromo-electric field energy contribution to the Hamiltonian becomes
@@ III.1 Color Entanglement
In the case of U(1) gauge theory, heavy charges can be consistently included by discontinuities in Gauss’s law and in values of the gauge field at each spatial location. This is implemented by mapping the field to qubits (the occupation basis), where each basis state is an eigenstate of the electric charge operator. [^2] When acted on by the electric charge operator, an electron remains an electron, and a positron remains a positron. In 1+1D QED, external charges are screened and charge-charge correlation functions fall off exponentially outside of a few confinement lengths. The situation is more complex in the case of non-Abelian theories, even in 1+1D. Specifically, because the non-Abelian charge operators act non-trivially on the fields, specifically, mixing states in the occupation basis, external charges cannot be efficiently included via discontinuities in Gauss’s law and in the chromoelectric fields (see related works using tensor networks [Kuhn:2015zqa, Sala:2018dui]). [^3] This leads to the use of a quadratic penalty for the energy in the electric field beyond the end of the lattice to enforce global color neutrality. Hence, heavy quarks are mapped to qubits in the same way as the light quarks, but a kinetic term is not included in the Hamiltonian to prevent its delocalization from a given lattice site.
@@
With a heavy-quark present, its non-Abelian charge can be screened in two ways. The first is by the presence of light anti-quarks, corresponding to the presence of (color-singlet) heavy mesons (with interpolating operators of the form $\hat{O}\sim\sum_{\alpha}\overline{Q}{}^{\alpha}q_{\alpha}$). The second is by the presence of light quarks, corresponding to the presence of (color-singlet) heavy baryons (with interpolating operators of the form $\hat{O}\sim\sum_{\alpha\beta}\epsilon_{\alpha\beta}\ \overline{Q}{}^{\alpha}\overline{q}{}^{\beta}$). [^5] The Abelian theory does not have corresponding baryon operators. The latter is energetically favored (in the lattice theory) as there is no chromo-electric flux between the quark and anti-quark sectors. [^6] As noted in Ref. [Atas:2021ext], the baryon and meson masses converge in the continuum limit where lattice-spacing artifacts vanish. The SC ground-state wavefunction for this $n_{Q}=1$ system is of the form [^7] The (entangled) wavefunction created from a meson interpolating operator is of the form $|\ Q_{r}\ Q_{g}\ q_{r}\ q_{g}\ \overline{q}_{r}\ \overline{q}_{g}\ \rangle\rightarrow\frac{1}{\sqrt{2}}\biggl[|\uparrow\downarrow\downarrow\downarrow\ \downarrow\uparrow\rangle\ +\ |\downarrow\uparrow\downarrow\downarrow\ \uparrow\downarrow\rangle\biggl]\ =\ \frac{1}{\sqrt{2}}\biggl[|011110\rangle\ +\ |101101\rangle\biggr]\ .$ (10)
@@ III.2 Implementing Heavy-Quark Motion
For simulations of the Schwinger model, continuous spatial translations of the classical background charge can be implemented by a position-dependent reweighting of the discontinuities in the Gauss’s law constraint [Florio:2023dke, Florio:2024aix, Farrell:2024mgu]. For the reasons mentioned above, specifically the action of the non-Abelian charge operator, spatial translations in non-Abelian theories cannot be efficiently used in this way, and instead are implemented by fermionic-SWAP (FSWAP) operators [Verstraete:2008qpa, Cervera-Lierta:2018bia, Algaba:2023enr] acting among the heavy-quark lattice sites. Acting on a two-qubit system $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$, the continuous FSWAP operation is
@@ III.3 Ground-State Wavefunctions with Sources, and Domain Decomposition
In previous works, we developed scalable quantum circuits [Farrell:2023fgd, Farrell:2024fit] implemented by ADAPT-VQE [Grimsley_2019], called SC-ADAPT-VQE (for a variant, see Ref. [Gustafson:2024bww]). Classical computing was used to determine the operator sequence and their coefficients in creating a state approaching the ground state in modest-sized volumes, and then scaling to arbitrarily large systems on quantum computers in a controlled and converging way using the symmetries of the (lattice) system. When wavepackets or heavy-quark sources are included, localized regions of the system are modified from the vacuum by operators with finite extent to approach the target wavefunction.
@@ IV.2 Hadronic Interpolating Operators for State Preparation
$\hat{O}_{M1}\equiv(\hat{Y}\hat{Z}^{3}\hat{X}-\hat{X}\hat{Z}^{3}\hat{Y})\hat{I}+\hat{I}(\hat{Y}\hat{Z}^{3}\hat{X}-\hat{X}\hat{Z}^{3}\hat{Y})\ ,\ \hat{O}_{B1}\equiv-4i(\hat{\sigma}^{+}\hat{\sigma}^{+}\hat{Z}^{2}\hat{\sigma}^{-}\hat{\sigma}^{-}-\rm{h.c.})\ ,$ (33)
where, as for the $L=1$ system, the $\hat{X},\hat{Y}$ operators only act on the light quarks and anti-quarks. These are both operators that “straddle” heavy $Q$ sites, i.e., the extra $\hat{Z}^{2}$ acts on $Q_{r},Q_{g}$ sites. A more detailed presentation of this operator pool can be found in Eq. (83). In general, for arbitrary systems, meson and baryon-antibaryon operators, or in fact any color-singlet configurations, with support across multiple sites, located at arbitrary positions on the lattice should be included. Their relative contributions are determined by their (hadronic) form factors and the structure of the ground-state wavefunction.
@@
Figure 3 shows the convergence of the $\langle\hat{Z}_{j}\rangle$ as a function of the fermion site $j$, starting from $|SC;2,0\rangle$ in the $n_{Q}=0$ sector. The small icons added to the figures throughout this manuscript represent wether they have been obtained using a classical simulator (purple triangle) or with quantum hardware (blue diamond) [Klco:2019xro]. The meson layers are seen to converge the (single qubit) $\langle\hat{Z}_{j}\rangle$ to be close to their exact values, but the infidelity of the wavefunction is above $\sim 2\%$. Including the baryon-antibaryon operators reduces the infidelity to below $\sim 0.1\%$.

### v3

@@ I Introduction
Extensive experimental campaigns in high-energy heavy-ion collisions are designed to reveal the structure and dynamics of matter under extreme conditions. The remarkable discovery of a near-ideal, low-viscosity quantum liquid of quarks and gluons produced in such collisions [1, 2, 3, 4] has driven a program to measure and interpret its properties. An important line of investigation is to quantify properties of transport in the liquid phase, where the identification of signatures and probes present a challenge. One class of probes is the production, motion and measurement of heavy quarks, including the production rates of quarkonia, and processes such as dissociation and recombination. Produced in hard collisions at early times, heavy quarks witness the entire evolution of the exotic matter and its transition to hadronic final states. Therefore, the measured relative yields of heavy-quark systems reveal key attributes of the medium. An excellent review of the state-of-the-art in this area can be found in Ref. [5]. An important element in extracting properties of this exotic matter is understanding how heavy quarks move and interact within the evolving system, including how they lose and gain energy as they move. From a theoretical standpoint, first-principles simulations of these collisions present a monumental challenge. This challenge is not limited to evolution through exotic matter. Even simulations of high-energy probes impacting cold dense matter, such as a nucleus, or nuclear matter, suffer the same difficulties. While processes occurring over short-distances can be reliably determined in perturbation theory, the subsequent evolution (cascade) to longer length scales involves fragmentation and hadronization, both of which are intrinsically non-perturbative. Tuned to experimental data, modeling such processes is impressively successful, e.g., Geant4 [6, 7, 8, 9], but remains incomplete in describing fundamentally quantum observables, e.g., Refs. [10, 11, 12, 13, 14, 15]. While classical Euclidean-space simulations of quantum chromodynamics (QCD) encounter a sign problem in summing complex phases in the time-evolution, real-time evolution, starting from a specified initial state, is known to be efficient for quantum computers.
@@
Constrained by available noisy intermediate-scale quantum (NISQ) [16] era quantum hardware, present-day quantum simulations of quantum field theories are focused on modest-sized, low-dimensional systems [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73], with the Schwinger model [74] (quantum electrodynamics (QED) in 1+1D) receiving significant attention. Like QCD, the Schwinger model is a confining gauge theory, and remains a useful test-bed for exploring new algorithms and workflows and gaining improved understanding of infinite-volume and continuum limits. Our present work is related to simulating transport, in a general sense, in non-equilibrium non-Abelian systems, and is on the path for robust simulations of jet physics and hadronization in QCD. There is a growing foundation for such simulations, e.g., Refs. [75, 76, 77, 78, 79, 80, 81, 25, 82, 83, 84, 85, 86, 87, 88, 89, 90, 65, 91], along with progress in string breaking [92, 93, 94, 95, 96, 97, 62, 59, 98, 99, 100, 101, 102], and of the associated entanglement evolution [103, 104]. Simulations of SU(2) and SU(3) theories including quarks and anti-quarks in 1+1D, with the gluon fields constrained by Gauss’s law (either as dynamical fields or through long-range interactions), are now being performed using quantum computers [21, 24, 26, 40, 42, 41, 63, 69]. With the Jordan-Wigner mapping of quarks, each spatial lattice site requires $2n_{f}N_{c}$ qubits for $n_{f}$ flavors of light quarks with $N_{c}$ colors. The inclusion of electroweak processes [42, 69] (using the low-energy effective field theory of the Standard Model) requires additional $4n_{g}$ qubits encoding the charged-leptons and neutrinos, where $n_{g}$ is the number of generations included in the simulation. Progress has also been made toward simulations using qudit systems, e.g., Refs. [105, 49, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117], which have significant advantages in circuit depth and execution times. Additionally, there have been a number of important quantum and classical algorithm developments for wavepacket preparation for scattering [103, 53, 118, 55, 60, 119, 66, 120, 67], and for systematically-improvable truncation schemes [121].
@@
Building on our previous work that established a framework for quantum simulations of energy-loss and fragmentation in systems of electrically-charged particles in the context of Abelian lattice gauge theory (LGT) [118], we now address energy loss in non-Abelian lattice gauge theories. [^1] This should be considered to be a specific case of quantum thermodynamics of a non-equilibrium process [122, 123], in the context of a non-Abelian lattice gauge theory. In particular, we establish a framework to simulate the energy loss and collisions of particles in an SU(2) non-Abelian lattice gauge theory in 1+1D using quantum computers. Static heavy quark fields are mapped to qubits in order to simulate hadron motion and collisions on the lattice (as well as to establish regions of dense matter). To prepare ground-state wavefunctions in the presence of a heavy quark, SC-ADAPT-VQE [124, 48] is combined with Domain Decomposition (DDec), an algorithm used extensively in traditional lattice QCD calculations [125, 126, 127, 128, 129, 130], leveraging the heterogeneity of charge sectors across different regions on the lattice. The heavy quarks are translated across the lattice via discrete fermionic-SWAP operators, introducing additional finite lattice-spacing artifacts that vanish in the continuum limit. The states are then time-evolved to allow the light quarks to respond dynamically to the heavy quark motion. Scalable quantum circuits [48] are developed for state preparation, discrete translation, and time evolution, enabling measurements of the time dependence of local non-Abelian charge fluctuations, entanglement, and the energy in the light quarks and gluons. Complete simulations (state preparation, translation, time evolution and measurement, with associated mitigation circuits) of a lattice with $L=3$ spatial sites are performed, requiring a total of 18 qubits, using IBM’s heron-processor quantum computers, with a transpiled CNOT depth of $\sim 400$. With multiple error-mitigation strategies applied, the results we obtain from the quantum simulations are in good agreement with corresponding classical simulations.
@@ II The SU(2) Lattice with Heavy Quarks
Contributions to the Hamiltonian from the chromo-electric fields are determined by combinations of the non-Abelian SU(2) charge operators [26],
@@
where $\lfloor y\rfloor$ denotes the floor of $y$. With OBCs, and the requirement that the chromo-electric field outside of the lattice ($x<0$ and x>L1x>L-1) has to vanish, the summation is restricted to the electric-field boundaries in the interior of the lattice, $b\in\{0,1,\ldots,2L-2\}$. Color neutrality of the lattice is enforced by including a large energy penalty term in the Hamiltonian for non-zero $E_{b}^{(a)}$ at b=2L1b=2L-1 [40]. The chromo-electric field energy contribution to the Hamiltonian becomes
@@ III.1 Color Entanglement
In the case of U(1) gauge theory, heavy charges can be consistently included by discontinuities in Gauss’s law and in values of the gauge field at each spatial location. This is implemented by mapping the field to qubits (the occupation basis), where each basis state is an eigenstate of the electric charge operator. [^2] When acted on by the electric charge operator, an electron remains an electron, and a positron remains a positron. In 1+1D QED, external charges are screened and charge-charge correlation functions fall off exponentially outside of a few confinement lengths. The situation is more complex in the case of non-Abelian theories, even in 1+1D. Specifically, because the non-Abelian charge operators act non-trivially on the fields, specifically, mixing states in the occupation basis, external charges cannot be efficiently included via discontinuities in Gauss’s law and in the chromoelectric fields (see related works using tensor networks [75, 77]). [^3] This leads to the use of a quadratic penalty for the energy in the electric field beyond the end of the lattice to enforce global color neutrality. Hence, heavy quarks are mapped to qubits in the same way as the light quarks, but a kinetic term is not included in the Hamiltonian to prevent its delocalization from a given lattice site.
@@
With a heavy-quark present, its non-Abelian charge can be screened in two ways. The first is by the presence of light anti-quarks, corresponding to the presence of (color-singlet) heavy mesons (with interpolating operators of the form $\hat{O}\sim\sum_{\alpha}\overline{Q}{}^{\alpha}q_{\alpha}$). The second is by the presence of light quarks, corresponding to the presence of (color-singlet) heavy baryons (with interpolating operators of the form $\hat{O}\sim\sum_{\alpha\beta}\epsilon_{\alpha\beta}\ \overline{Q}{}^{\alpha}\overline{q}{}^{\beta}$). [^5] The Abelian theory does not have corresponding baryon operators. The latter is energetically favored (in the lattice theory) as there is no chromo-electric flux between the quark and anti-quark sectors. [^6] As noted in Ref. [26], the baryon and meson masses converge in the continuum limit where lattice-spacing artifacts vanish. The SC ground-state wavefunction for this $n_{Q}=1$ system is of the form [^7] The (entangled) wavefunction created from a meson interpolating operator is of the form $|\ Q_{r}\ Q_{g}\ q_{r}\ q_{g}\ \overline{q}_{r}\ \overline{q}_{g}\ \rangle\rightarrow\frac{1}{\sqrt{2}}\biggl[|\uparrow\downarrow\downarrow\downarrow\ \downarrow\uparrow\rangle\ +\ |\downarrow\uparrow\downarrow\downarrow\ \uparrow\downarrow\rangle\biggl]\ =\ \frac{1}{\sqrt{2}}\biggl[|011110\rangle\ +\ |101101\rangle\biggr]\ .$ (10)
@@ III.2 Implementing Heavy-Quark Motion
For simulations of the Schwinger model, continuous spatial translations of the classical background charge can be implemented by a position-dependent reweighting of the discontinuities in the Gauss’s law constraint [88, 90, 118]. For the reasons mentioned above, specifically the action of the non-Abelian charge operator, spatial translations in non-Abelian theories cannot be efficiently used in this way, and instead are implemented by fermionic-SWAP (FSWAP) operators [131, 132, 133] acting among the heavy-quark lattice sites. Acting on a two-qubit system $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$, the continuous FSWAP operation is
@@ III.3 Ground-State Wavefunctions with Sources, and Domain Decomposition
In previous works, we developed scalable quantum circuits [48, 53] implemented by ADAPT-VQE [124], called SC-ADAPT-VQE (for a variant, see Ref. [134]). Classical computing was used to determine the operator sequence and their coefficients in creating a state approaching the ground state in modest-sized volumes, and then scaling to arbitrarily large systems on quantum computers in a controlled and converging way using the symmetries of the (lattice) system. When wavepackets or heavy-quark sources are included, localized regions of the system are modified from the vacuum by operators with finite extent to approach the target wavefunction.
@@ IV.2 Hadronic Interpolating Operators for State Preparation
$\hat{O}_{M1}\equiv(\hat{Y}\hat{Z}^{3}\hat{X}-\hat{X}\hat{Z}^{3}\hat{Y})\hat{I}+\hat{I}(\hat{Y}\hat{Z}^{3}\hat{X}-\hat{X}\hat{Z}^{3}\hat{Y})\ ,\ \hat{O}_{B1}\equiv-4i(\hat{\sigma}^{+}\hat{\sigma}^{+}\hat{I}^{2}\hat{\sigma}^{-}\hat{\sigma}^{-}-\rm{h.c.})\ ,$ (33)
where, as for the $L=1$ system, the $\hat{X},\hat{Y}$ operators only act on the light quarks and anti-quarks. These are both operators that “straddle” heavy $Q$ sites, i.e., the extra $\hat{Z}^{2}$ or $\hat{I}^{2}$ acts on $Q_{r},Q_{g}$ sites. A more detailed presentation of this operator pool can be found in Eq. (83). In general, for arbitrary systems, meson and baryon-antibaryon operators, or in fact any color-singlet configurations, with support across multiple sites, located at arbitrary positions on the lattice should be included. Their relative contributions are determined by their (hadronic) form factors and the structure of the ground-state wavefunction.
@@
Figure 3 shows the convergence of the $\langle\hat{Z}_{j}\rangle$ as a function of the fermion site $j$, starting from $|SC;2,0\rangle$ in the $n_{Q}=0$ sector. The small icons added to the figures throughout this manuscript represent wether they have been obtained using a classical simulator (purple triangle) or with quantum hardware (blue diamond) [135]. The meson layers are seen to converge the (single qubit) $\langle\hat{Z}_{j}\rangle$ to be close to their exact values, but the infidelity of the wavefunction is above $\sim 2\%$. Including the baryon-antibaryon operators reduces the infidelity to below $\sim 0.1\%$.